A block of wood weigh 4N is placed on a horizontal table it is then pulled by means of a stringe if the tension in the stringe is 2.5N calculate the coefficient of the static friction
1
Expert's answer
2021-10-21T10:52:38-0400

The tension in the string equals in magnitude to the force of friction. Thus, the coefficient of friction will be


"\mu=\frac fN=\frac fW=\frac TW=0.625."